Stevie is twice as old as his brother. Four years ago he was four times as old a s his brother. Find their present ages.
Please help with word problem, I can't come up with an equation.

Answer by jonvaliente(64)

Let x=Steve's borther's age
2x=Steve's age
x-4 = 4*(2x-4)
x-4 = 8x-16
7x=12
x=12/7 brother's age
2x=24/7 steve's age
There seems to be something missing in the problem.
Obviously, both brothers weren't even born 4 years ago so this can't be right.
But I believe the math is right for this one though.